# Curvature pinching of asymptotically conical gradient expanding Ricci solitons

### v2

@@ 3. Curvature pinching for asymptotically conical Ricci expanders with vanishing Weyl tensor
$\bullet$ Subcase 1a: $\lambda_{2}=0$ at $p_{0}$. By Theorem 3.1 (i), we have $\operatorname{Ric}\geq 0$, which implies $\lambda_{1}(p_{0})=0$. Then, the third relation in (3.10) yields $\lambda_{i}=\lambda_{j}$ at $p_{0}$ for all $i,j\geq 2$, hence $\lambda_{i}(p_{0})=0$ for all $i$, contradicting the assumption that the scalar curvature $R$ is positive.
@@
Case 2: Infimum at infinity. Suppose $\mu/R$ attains its negative infimum at infinity. Then, by Lemma 6.2 and essentially the same argument as in the proof of $Rc\geq 0$ in Part I, we can rule out Case 2. We omit the details.
Hence, we conclude that $M^{-}=\emptyset$. Therefore, $M$ has nonnegative sectional curvature. As nonnegative sectional curvature is equivalent to nonnegative curvature operator for manifolds with vanishing Weyl tensor $W=0$, it follows that the expanding soliton $(M,g,f)$ has nonnegative curvature operator $Rm\geq 0$. In fact, $\mu$ coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [89, Proposition 3.1(i)]).
Finally, it remains to prove $(M,g,f)$ has positive curvature operator $Rm>0$. To begin with, following the proof of [25, Theorem 1.3], we observe that the null space of the curvature operator, $\ker(Rm)$, is invariant under parallel translation. Indeed, consider the canonical immortal solution
@@
and the fact that $(M,g)$ has nonnegative curvature operator $Rm\geq 0$, it follows that the quadratic term $Rm(t)^{2}+Rm(t)^{\sharp}\geq 0$ for all $t\in[0,1]$. Thus, by Hamilton’s strong maximum principle (see, e.g., [27, Theorem 2.2.1]), there exists an interval $0<t<\delta$ over which the rank of $Rm(t)$ is constant, and $\ker(Rm(t))$ is invariant under parallel translation. Since $g(t)=(1+t)\Phi(t)^{\ast}(g)$ with $g(0)=g$, we conclude that the rank of $Rm=Rm(0)$ is locally constant and that $\ker(Rm)=\ker(Rm(0))$ is invariant under parallel translation.
@@
Proof of Claim. As in[87], we argue by contradiction. Suppose that $M$ does not have positive curvature operator. Since $M$ has nonnegative curvature operator, there exists a point $p_{0}\in M$ such that the smallest eigenvalue of the curvature operator $\mu(p_{0})=0$. Because the rank of the curvature operator $Rm$ is locally constant, it follows that $\mu\equiv 0$ on a neighborhood $U$ of $p_{0}$. Moreover, as $\ker(Rm)$ is invariant under parallel translation, there exists an orthogonal decomposition of the tangent bundle $TU=V_{1}\oplus V_{2}$, where $V_{1}=\ker(Rm)$ with $\dim(V_{1})\geq 1$, and both $V_{1}$ and $V_{2}$ are invariant under parallel translation. By [60, Lemma 9.1], $M$ locally splits off at least one Euclidean factor. Since expanding Ricci solitons are real analytic, completeness implies that this local splitting extends globally. However, this contradicts the assumption that $M$ is asymptotically conical: any Euclidean factor of $M$ would force the asymptotic cone $\mathcal{C}$ to split off a Euclidean factor, thereby violating the regularity assumption that $\mathcal{C}$ has only an isolated singularity at the tip, unless $\mathcal{C}$ is a Euclidean space with flat metric. In this latter case, a flat Euclidean space is not a cone with positive scalar curvature, which is a contradiction. Consequently, $M$ must have positive curvature operator.
@@
Let $(M^{4},g,f)$ be a non-compact, half conformally flat (i.e., either $W^{+}=0$ or $W^{-}=0$), asymptotically conical Ricci expander with positive scalar curvature. By essentially the same arguments as in Chen-Wang [39] for gradient shrinking and steady Ricci solitons, one can show that $(M^{4},g,f)$ has vanishing $D$-tensor as introduced in [18, 19]. Thus, by [19, Proposition 3.2], $(M^{4},g,f)$ is locally conformally flat. Therefore, it follows from Theorem 3.1 that $(M^{4},g,f)$ has positive curvature operator $Rm>0$. ∎
Finally, we conclude this section by proving Theorem 1.1.
@@
First, note that we can diagonalize the curvature operator $Rm$ with eigenvalues $m_{1}\leq m_{2}\leq m_{3}$. Then, the Ricci tensor $Rc$ is diagonalized with eigenvalues
@@
Case 2: Infimum at infinity. Assume instead that $m_{1}/R$ attains its negative infimum at infinity. By Lemma 6.1, any 3-dimensional cone with positive scalar curvature automatically has positive curvature operator/sectional curvature, and the smallest eigenvalue of its curvature operator is zero. Then, as in Part I of the proof of Theorem 3.1, we get a contradiction. Thus, Case 2 is impossible.
@@
Finally, by the same argument as in the proof of $Rm>0$ in Theorem 3.1, we conclude that $M^{3}$ has positive curvature operator. This completes the proof of Theorem 1.1. ∎
Remark 3.2.
Theorem 1.1 also follows from Theorem 1.2, which is valid for $n=3$ since the Weyl tensor vanishes automatically, and Lemma 6.1 on the curvature operator of 3D cones.
@@
(b) Without loss of generality, we assume $A_{2}>0$. By part (a), this implies $A_{1}\geq 0$. Following [24, Proposition 3.1(b)], we prove that $A_{1}>0$ by contradiction. Suppose $A_{1}(p_{0})=0$ at some point $p_{0}\in M^{4}$. Then $A_{1}$ attains its minimum at $p_{0}$. Let $\eta\in\wedge_{p_{0}}^{+}(M)$ be a null eigenvector of $A$ such that $A(\eta,\eta)=A_{1}(p_{0})=0$ at $p_{0}$. Extend $\eta$ to a local section (still denoted by $\eta$) by parallel transport along geodesics emanating from $p_{0}$.
@@
Let $0\leq B_{1}\leq B_{2}\leq B_{3}$ be the singular eigenvalues of the matrix $B$ and ${\lambda}_{1}\leq{\lambda}_{2}\leq{\lambda}_{3}\leq{\lambda}_{4}$ be the eigenvalues of the Ricci tensor Rc{Rc}. Then, by [25, Lemma 2.2], the sum of the least two eigenvalues of $Rc$ is given by
@@
Now, by Lemma 2.1 and essentially the same computations as in the proof of [25, Theorem 3.1(a)], we have
@@
$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{2}{R^{2}}\big[u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\big],\end{split}$ (4.1)
@@
which implies $u|Rc|^{2}=4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}=0$ on $\Omega$.
@@
Case 2: Infimum at infinity. Suppose that $u/R$ attains its negative infimum at infinity. Since $(M^{4},g,f)$ has PIC, the asymptotic cone must have WPIC. By Lemma 6.3, the Ricci curvature of the asymptotic cone is therefore nonnegative. Then by similar arguments in the proof of Theorem 3.1, we can rule out Case 2.
@@
Finally, given that $(M^{4},g,f)$ has 2-nonnegative Ricci curvature, following the proof in [25, Theorem 3.1(b)], we shall prove 2-positive Ricci curvature by contradiction. We consider the quadratic form $Z:=RI-4\sqrt{B{}^{t}B}$, where $I$ is the 3 by 3 identity matrix. By the 2-nonnegativity of the Ricci curvature, we know that $Z\geq 0$ and that 2-positive Ricci curvature is equivalent to $Z>0$. Now, we denote the eigenvalues of $Z$ by
@@
This completes the proof of Theorem 1.4. ∎
Finally, by Lemma 6.3 and an argument similar to that used in the proof of Theorem 1.4, we obtain the following slightly stronger result than Theorem 1.4.
@@
    If $(M^{4},g,f)$ satisfies $A_{2}\geq 0$ and $C_{2}\geq 0$, then the Ricci curvature is $2$-nonnegative.
@@
    If $(M^{4},g,f)$ satisfies $A_{2}>0$ and $C_{2}>0$, then the Ricci curvature is $2$-positive.
@@ 5. General lemma and further applications
In this section, we formulate a fairly general method that can be effectively applied to prove generalized Hamilton-Ivey type curvature pinching estimates for a class of non-compact, asymptotically conical gradient Ricci solitons. As applications, we obtain several analogues of other known curvature pinching results for ancient solutions, in the setting of asymptotically conical expanding Ricci solitons–including Theorem 1.5 as stated in the introduction, and Theorem 5.1 below.
@@ 5.1. A general lemma
Let $(M^{n},g,f)$ be an $n$-dimensional non-compact gradient expanding Ricci soliton, satisfying Eq. (1.1) and with positive scalar curvature. Suppose $u:M^{n}\rightarrow\mathbb{R}$ is a Lipschitz function and satisfies the differential inequality
@@
$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&\leq\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{1}{R}(2\rho u-v)-\frac{u}{R^{2}}(2\rho R-2|Rc|^{2})\\ &\leq\frac{1}{R^{2}}(2u|Rc|^{2}-Rv).\end{split}$ (5.4)
@@
Lemma 5.1 also applies to gradient shrinking and steady Ricci solitons.
@@
First of all, it is easy to see that the inequality $B_{3}^{2}\leq A_{1}C_{1}$ implies nonnegative curvature operator $Rm\geq 0$ for $(M^{4},g)$; see, e.g., [43, Lemma 4.4] for a proof. Hence, to prove Theorem 1.5, it suffices to establish the inequality $B_{3}^{2}\leq A_{1}C_{1}$.
@@
By the same computations as in [7, Lemma 6.1], we have
@@
Moreover, since the asymptotic cone $\mathcal{C}$ is a non-flat Euclidean cone, by (6.8) and (6.9), on the cone $\mathcal{C}$, we have
@@
Let $w_{1}:=\tfrac{1}{2}\sqrt{(A_{1}+A_{2})(C_{1}+C_{2})}$. By the same computations as in [7, Lemma 6.2], we can find a positive constant $\delta_{1}>0$ such that
@@
Let $w_{2}:=\sqrt{A_{1}C_{1}}$. Then by the same computations as in [7, Proposition 6.3], we can find a positive constant $\delta_{2}>0$ such that
@@
The assumption of the asymptotic cone being a (non-flat) Euclidean cone is used in the proofs of Step 2 and Step 3. Indeed, by (6.8) and (6.9), requiring either $4B_{3}^{2}\leq(A_{1}+A_{2})(C_{1}+C_{2})$ or $B_{3}^{2}\leq A_{1}C_{1}$ to hold on the asymptotic cone forces the link of the cone to be a (spherical) space form.

### v3

@@ 3. Curvature pinching for asymptotically conical Ricci expanders with vanishing Weyl tensor
$\bullet$ Subcase 1a: $\lambda_{2}=0$ at $p_{0}$. By Part I, we have $\operatorname{Ric}\geq 0$, which implies $\lambda_{1}(p_{0})=0$. Then, the third relation in (3.10) yields $\lambda_{i}=\lambda_{j}$ at $p_{0}$ for all $i,j\geq 2$, hence $\lambda_{i}(p_{0})=0$ for all $i$, contradicting the fact that the scalar curvature $R$ is positive.
@@
Case 2: Infimum at infinity. Suppose $\mu/R$ attains its negative infimum at infinity. Then, by Lemma A.1 and Lemma A.2 and essentially the same argument as in the proof of $Rc\geq 0$ in Part I, we can rule out Case 2. We omit the details.
Hence, we conclude that $M^{-}=\emptyset$. Therefore, $M$ has nonnegative sectional curvature. As nonnegative sectional curvature is equivalent to nonnegative curvature operator for manifolds with vanishing Weyl tensor $W=0$, it follows that the expanding soliton $(M,g,f)$ has nonnegative curvature operator $Rm\geq 0$. In fact, $\mu$ coincides with the smallest eigenvalue of the curvature operator $Rm$ (see, e.g., [90, Proposition 3.1(i)]). ∎
Part III: $(M^{n},g,f)$ has positive curvature operator $Rm>0$.
Proof of $Rm>0$.
To begin with, following the proof of [27, Theorem 1.3], we observe that the null space of the curvature operator, $\ker(Rm)$, is invariant under parallel translation. Indeed, consider the canonical immortal solution
@@
and the fact that $(M,g)$ has nonnegative curvature operator $Rm\geq 0$, it follows that the quadratic term $Rm(t)^{2}+Rm(t)^{\sharp}\geq 0$ for all $t\in[0,1]$. Thus, by Hamilton’s strong maximum principle (see, e.g., [29, Theorem 2.2.1]), there exists an interval $0<t<\delta$ over which the rank of $Rm(t)$ is constant, and $\ker(Rm(t))$ is invariant under parallel translation. Since $g(t)=(1+t)\Phi(t)^{\ast}(g)$ with $g(0)=g$, we conclude that the rank of $Rm=Rm(0)$ is locally constant and that $\ker(Rm)=\ker(Rm(0))$ is invariant under parallel translation.
@@
Proof of Claim. As in [88], we argue by contradiction. Suppose that $M$ does not have positive curvature operator. Since $M$ has nonnegative curvature operator, there exists a point $p_{0}\in M$ such that the smallest eigenvalue of the curvature operator $\mu(p_{0})=0$. Because the rank of the curvature operator $Rm$ is locally constant, it follows that $\mu\equiv 0$ on a neighborhood $U$ of $p_{0}$. Moreover, as $\ker(Rm)$ is invariant under parallel translation, there exists an orthogonal decomposition of the tangent bundle $TU=V_{1}\oplus V_{2}$, where $V_{1}=\ker(Rm)$ with $\dim(V_{1})\geq 1$, and both $V_{1}$ and $V_{2}$ are invariant under parallel translation. By [61, Lemma 9.1], $M$ locally splits off at least one Euclidean factor. Since expanding Ricci solitons are real analytic, completeness implies that this local splitting extends globally. However, this contradicts the assumption that $M$ is asymptotically conical: any Euclidean factor of $M$ would force the asymptotic cone $\mathcal{C}$ to split off a Euclidean factor, thereby violating the regularity assumption that $\mathcal{C}$ has only an isolated singularity at the tip, unless $\mathcal{C}$ is a Euclidean space with flat metric. In this latter case, a flat Euclidean space is not a cone with positive scalar curvature, which is a contradiction. Consequently, $M$ must have positive curvature operator.
@@
Let $(M^{4},g,f)$ be a non-compact, half-conformally flat (i.e., either $W^{+}=0$ or $W^{-}=0$), asymptotically conical Ricci expander with positive scalar curvature. By essentially the same arguments as in Chen-Wang [41] for gradient shrinking and steady Ricci solitons, one can show that $(M^{4},g,f)$ has vanishing $D$-tensor as introduced in [20, 21]. Thus, by [21, Proposition 3.2], $(M^{4},g,f)$ is locally conformally flat. Therefore, it follows from Theorem 3.1 that $(M^{4},g,f)$ has positive curvature operator $Rm>0$. ∎
Finally, we conclude this section by providing a more direct proof of Theorem 1.1.
@@
First, note that we can diagonalize the curvature operator $Rm$ with eigenvalues $m_{1}\leq m_{2}\leq m_{3}$, which are principal sectional curvatures of $(M^{3},g,f)$, with respect to a suitable orthonormal frame $\{e_{1},e_{2},e_{3}\}$ at each point $p\in M^{3}$. Then, the Ricci tensor $Rc$ is diagonalized with eigenvalues
@@
Case 2: Infimum at infinity. Assume instead that $m_{1}/R$ attains its negative infimum at infinity. By Lemma A.1, any 3-dimensional cone with positive scalar curvature automatically has positive curvature operator/sectional curvature, and the smallest eigenvalue of its curvature operator is zero. Then, as in Part I of the proof of Theorem 3.1, we get a contradiction. Thus, Case 2 is impossible.
@@
Finally, by the same argument as in the proof of $Rm>0$ (i.e., Part III) for Theorem 3.1, we conclude that $M^{3}$ has positive curvature operator. This completes the proof of Theorem 1.1. ∎
@@
(b) Without loss of generality, we assume $A_{2}>0$. By part (a), this implies $A_{1}\geq 0$. Following [26, Proposition 3.1(b)], we prove that $A_{1}>0$ by contradiction. Suppose $A_{1}(p_{0})=0$ at some point $p_{0}\in M^{4}$. Then $A_{1}$ attains its minimum at $p_{0}$. Let $\eta\in\wedge_{p_{0}}^{+}(M)$ be a null eigenvector of $A$ such that $A(\eta,\eta)=A_{1}(p_{0})=0$ at $p_{0}$. Extend $\eta$ to a local section (still denoted by $\eta$) by parallel transport along geodesics emanating from $p_{0}$.
@@
Let $0\leq B_{1}\leq B_{2}\leq B_{3}$ be the singular eigenvalues of the matrix $B$ and ${\lambda}_{1}\leq{\lambda}_{2}\leq{\lambda}_{3}\leq{\lambda}_{4}$ be the eigenvalues of the Ricci tensor Rc{Rc}. Then, by [27, Lemma 2.2], the sum of the least two eigenvalues of $Rc$ is given by
@@
Now, by Lemma 2.1 and essentially the same computations as in the proof of [27, Theorem 3.1(a)], we have
@@
$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{2}{R^{2}}\big[u|Rc|^{2}-4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\big].\end{split}$ (4.1)
@@
which implies $u|Rc|^{2}=4R(A_{2}+A_{1}+C_{2}+C_{1})B_{3}\equiv 0$ on $\Omega$.
@@
Case 2: Infimum at infinity. Suppose that $u/R$ attains its negative infimum at infinity. Since $(M^{4},g,f)$ has PIC, the asymptotic cone must have WPIC. By Lemma A.3, the Ricci curvature of the asymptotic cone is therefore nonnegative. Then by similar arguments in the proof of Theorem 3.1, we can rule out Case 2.
@@
Given that $(M^{4},g,f)$ has 2-nonnegative Ricci curvature, following the proof in [27, Theorem 3.1(b)], we shall prove 2-positive Ricci curvature by contradiction. We consider the quadratic form $Z:=RI-4\sqrt{B{}^{t}B}$, where $I$ is the 3 by 3 identity matrix. By the 2-nonnegativity of the Ricci curvature, we know that $Z\geq 0$ and that 2-positive Ricci curvature is equivalent to $Z>0$. Now, we denote the eigenvalues of $Z$ by
@@
Finally, we establish the curvature estimate $|{Rm}|\leq 2R$. By the positive isotropic curvature condition, we have $|A|^{2}\leq\tfrac{3}{16}R^{2}$ and $|C|^{2}\leq\tfrac{3}{16}R^{2}$; see, e.g., the proof of [26, Theorem 1.3]. Since the Ricci tensor of $(M^{4},g,f)$ is $2$-positive, it follows, as in the proof of [27, Theorem 3.1(a)], that $|{Rc}|^{2}\leq R^{2}$. On the other hand,
$R^{2}\geq|{Rc}|^{2}=|\mathring{Rc}|^{2}+\tfrac{1}{4}R^{2}=4|B|^{2}+\tfrac{1}{4}R^{2}.$
Therefore,
$\begin{split}|{Rm}|^{2}&\leq 2\bigl(|A|^{2}+|C|^{2}+|B|^{2}\bigr)\leq\tfrac{9}{8}R^{2}.\end{split}$
This completes the proof of the curvature bound, and hence the proof of Theorem 1.4. ∎
Finally, by Lemma A.3 and an argument similar to that used in the proof of Theorem 1.4, we obtain the following slightly stronger result than Theorem 1.4.
@@
    If $(M^{4},g,f)$ satisfies $A_{2}\geq 0$ and $C_{2}\geq 0$, then the Ricci curvature is $2$-nonnegative, and $|Rm|\leq R$.
@@
    If $(M^{4},g,f)$ satisfies $A_{2}>0$ and $C_{2}>0$, then the Ricci curvature is $2$-positive, and $|Rm|\leq R$.
@@ 5. General lemma and further applications
In this section, we formulate a fairly general method that can be effectively applied to prove generalized Hamilton–Ivey type curvature pinching estimates for a class of non-compact, asymptotically conical gradient Ricci solitons. As applications, we obtain several analogues of other known curvature pinching results for ancient solutions, in the setting of asymptotically conical expanding Ricci solitons–including Theorem 1.5 as stated in the introduction, and Theorem 5.1 below.
@@ 5.1. A general lemma
Let $(M^{n},g,f)$ be an $n$-dimensional non-compact gradient Ricci soliton, satisfying Eq. (1.1) and with positive scalar curvature. Suppose $u:M^{n}\rightarrow\mathbb{R}$ is a Lipschitz function and satisfies the differential inequality
@@
$\begin{split}\Delta_{F}\left(\frac{u}{R}\right)&=\frac{1}{R}\Delta_{f}u-\frac{u}{R^{2}}\Delta_{f}R\\ &\leq\frac{1}{R}(2\rho u-v)-\frac{u}{R^{2}}(2\rho R-2|Rc|^{2})\\ &\leq\frac{1}{R^{2}}(2u|Rc|^{2}-Rv).\end{split}$ (5.4)
@@
Similarly, if a steady Ricci soliton $(M^{n},g,f)$ is asymptotic to a cylinder with $u\geq 0$, then $(M^{n},g,f)$ satisfies the asymptotic condition (5.2) in Lemma 5.1.
@@
First of all, it is easy to see that the inequality $B_{3}^{2}\leq A_{1}C_{1}$ implies nonnegative curvature operator $Rm\geq 0$ for $(M^{4},g)$; see, e.g., [44, Lemma 4.4] for a proof. Hence, to prove Theorem 1.5, it suffices to establish the inequality $B_{3}^{2}\leq A_{1}C_{1}$.
@@
By the same computations as in [8, Lemma 6.1], we have
@@
Moreover, since the asymptotic cone $\mathcal{C}$ is a non-flat Euclidean cone, by (A.9) and (A.10), on the cone $\mathcal{C}$, we have
@@
Let $w_{1}:=\tfrac{1}{2}\sqrt{(A_{1}+A_{2})(C_{1}+C_{2})}$. By the same computations as in [8, Lemma 6.2], we can find a positive constant $\delta_{1}>0$ such that
@@
Let $w_{2}:=\sqrt{A_{1}C_{1}}$. Then by the same computations as in [8, Proposition 6.3], we can find a positive constant $\delta_{2}>0$ such that
@@
The assumption of the asymptotic cone being a (non-flat) Euclidean cone is used in the proofs of Step 2 and Step 3. Indeed, by (A.9) and (A.10), requiring either $4B_{3}^{2}\leq(A_{1}+A_{2})(C_{1}+C_{2})$ or $B_{3}^{2}\leq A_{1}C_{1}$ to hold on the asymptotic cone forces the link of the cone to be a (spherical) space form.